One farmer says to another farmer: “Give me one of your cows, and then I will have twice as many cows as you have.” The second farmer replies: “Why don’t you just give me one of your cows, and then we will both have the same number of cows?” How many cows does each farmer have?
Answer by emathwiz(15)

Let the first farmer be x and second be y.
x+1 = 2(y-1)
y+1 = x-1
x = 2y-3
y+1= 2y-3-1
y=5 The second farmer has 5 cows
x+1=8
x=7 The first farmer has 7 cows